OpenStudy (jojokiw3):

My teacher says there's no solution to this question, but I'm getting answer and I figured that I probably missed something. Can someone check it?

OpenStudy (jojokiw3):

\[\log_{3} (x-2) - \log_{3} (x+1)= \log_{3} 5 \]

OpenStudy (jojokiw3):

I keep getting \[\frac{ -7 }{ 4 }\] as the answer for x.

OpenStudy (anonymous):

I'm getting the same.. and so does wolfram alpha http://www.wolframalpha.com/input/?i=log3%28x-2%29+-+log3%28x%2B1%29+%3D+log3%285%29 are you sure the question is right?

OpenStudy (jojokiw3):

Yeah I'm sure. I guess the teacher is probably wrong? o_O

OpenStudy (jojokiw3):

I looked at over 20 times lol.

OpenStudy (anonymous):

The thing is that wolfram writes '(assuming a complex-valued logarithm)' on the solution for this one, but not for others. let me think why it matters

OpenStudy (jojokiw3):

I don't even know what that is. -_-

OpenStudy (anonymous):

Ok, I think I got it. Think about it this way, is there a real power for 3 that will cause \(3^p \le 0\)? No, so n has to be positive in order for \(\log_3(n)\) to be able to return a real value. So if we say our log function has to work only with real numbers and cannot return complex numbers, then we can't take logarithm of non-positive numbers. So we have: $$ \log_3(x-2) - \log_3(x+1) = \log_3(5) $$That means we must have: $$ \begin{array}{ll|ll} x-2 > 0 &&& x+1 > 0\\ x > 2 &&& x > -1 \\ \hline \end{array}\\ \boxed{x>2} $$But our result is \(x = -\frac{7}{4}\) which is not in the valid range for \(x\), so there is no solution.

OpenStudy (jojokiw3):

Oh okay. Thanks chum!

OpenStudy (anonymous):

You're welcome =)